# Combinatorial flats and Schubert varieties of subspace arrangements

### v2

@@ 3. Combinatorial flats
By the matroid case and 2.15, a set $S$ is a flat of $\widetilde{T_{F}P}$ if and only if $S$ is a flat of $\widetilde{P}$ that either contains $\cup_{i\in F}E_{i}$ or satisfies $\overline{S\cup i}\not\supset\cup_{i\in F}E_{i}$ for all $i\in\widetilde{E}\setminus S$. In the former case, $\operatorname{rk}_{\widetilde{T_{F}P}}(S)=\operatorname{rk}_{\widetilde{P}}(S)-1$; in the latter $\operatorname{rk}_{\widetilde{T_{F}P}}(S)=\operatorname{rk}_{\widetilde{P}}(S)$. Taking $\mathbf{s}$ to be the multiset represented by $S$ completes the proof. ∎
In the remainder of Section 3, we work up to proving 1.1, then provide axioms for combinatorial flats. These results are not required to understand Schubert varieties of subspace arrangements; readers interested solely in geometry may proceed to Section 4 forthwith.
3.3.
@@
Semimodularity is 3.11.
For top-heaviness: let $H^{k}(\widetilde{P})$ be the $\mathbb{Q}$-vector space basis $\{v_{F}:F\in\mathcal{L}_{\widetilde{P}}^{k}\}$, and let $H^{k}(P,\mathbf{n})$ be the $\mathbb{Q}$-vector space with basis $\{v_{\mathbf{s}}:\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}^{k}\}$. From the quotient construction of $\mathcal{L}_{P,\mathbf{n}}$, there are isomorphisms
@@
where $S$ is any flat representing $\mathbf{s}$. By [BHM+20, Theorem 1.1], there is an injective linear $L:H^{k}(\widetilde{P})\to H^{d-k}(\widetilde{P})$ that commutes with the action of $\mathfrak{S}$. Consequently, the restriction of $L$ to $H^{k}(\widetilde{P})^{\mathfrak{S}}$ gives a linear injection
@@
and taking dimensions yields top-heaviness. Also by [BHM+20, Theorem 1.1], there are $\mathfrak{S}$-equivariant linear injections $H^{0}(\widetilde{P})\hookrightarrow H^{1}(\widetilde{P})\hookrightarrow\ldots\hookrightarrow H^{\lfloor d/2\rfloor}(\widetilde{P})$, which restrict to linear injections
@@
Generalizations of Lemmas 3.11 and 3.12 for matricubes appear in [AG24, Theorem 3.4] (see 3.8). With Lemmas 3.11 and 3.12 in hand, we have established 1.1 parts (i), (iii), and (iv). We now devote our attention to 1.1(ii), which states that $P$ is determined by $\mathcal{L}_{P}$ up to simplification.
@@
Since $P$ is simple, $E_{i}$ is a flat of $\widetilde{P}$ for all $i\in E$. The restriction of $\widetilde{P}$ to $E_{i}$ is a boolean matroid of rank $n_{i}=\operatorname{rk}_{P}(i)$, so all proposed join-irreducibles are combinatorial flats of $(P,\mathbf{n})$. Each $s_{i}\mathbf{e}_{i}$ is join-irreducible because the interval $[\mathbf{0},s_{i}\mathbf{e}_{i}]\subset\mathcal{L}_{P,\mathbf{n}}$ is a chain of length $s_{i}$. No other combinatorial flats are join-irreducible because $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$ can be written $\mathbf{s}=s_{1}\mathbf{e}_{1}\vee\cdots\vee s_{N}\mathbf{e}_{N}$. ∎
@@
By 3.13, each $i\in E$ corresponds to a maximal join-irreducible $x_{i}$ of $\mathcal{L}_{P,\mathbf{n}}$. The rank function of $P$ is recovered by $\operatorname{rk}_{P}(S)=\operatorname{rk}_{P,\mathbf{n}}(\vee_{i\in S}x_{i})$. ∎
@@
We show that reduction and loops do not change the lattice of combinatorial flats. From these facts, we derive 1.1.
@@
For surjectivity, suppose that $\mathbf{s}\in\mathcal{L}_{R_{i}(P,\mathbf{n})}$ is represented by a flat $S$ of $\widetilde{R_{i}(P,\mathbf{n})}$. We will show there is $\mathbf{s}^{\prime}\in\mathcal{L}_{P,\mathbf{n}}$ with $\psi(\mathbf{s}^{\prime})=\mathbf{s}$. By Lemmas 2.17 and 2.12, there is $e\in E_{i}\setminus S$ such that $S\cup e$ or $S$ is a flat of $\widetilde{P}$. If $S$ is a flat of $\widetilde{P}$, then $s_{i}<n_{i}$, so we may take $\mathbf{s}^{\prime}=\mathbf{s}$. Otherwise, suppose $S\cup e$ is a flat of $\widetilde{P}$. If $s_{i}+1<n_{i}$, then by 3.2, $\mathbf{s}$ is a combinatorial flat of $\widetilde{P}$, so we may again take $\mathbf{s}^{\prime}=\mathbf{s}$. Otherwise, if $S\cup e$ is a flat of $\widetilde{P}$ and $s_{i}+1=n_{i}$, then take $\mathbf{s}^{\prime}=\mathbf{s}+\mathbf{e}_{i}$. This completes the proof of surjectivity.
@@
Proof of 1.1.
Define $\mathcal{L}_{P}:=\mathcal{L}_{P,\mathbf{n}}$, with $\mathbf{n}$ any cage for $P$. This definition does not depend on the choice of cage because we may assume that $P$ is loopless by 3.16, then repeatedly apply 3.15 to produce an isomorphism $\mathcal{L}_{P,\mathbf{n}}\cong\mathcal{L}_{P,(\operatorname{rk}_{P}(1),\ldots,\operatorname{rk}_{P}(N))}$. We will henceforth assume that $\mathbf{n}=(\operatorname{rk}_{P}(1),\ldots,\operatorname{rk}_{P}(N))$.
Statements i and iv are the content of 3.12; we now move to proving ii. To construct $P^{\mathrm{sim}}$: let $(P^{\prime},\mathbf{n}^{\prime})=R_{i}(P\setminus L,\pi_{E\setminus L}(\mathbf{n}))$, where $L$ is the set of loops of $P$ and $i\in E\setminus L$ is any element such that $\{i\}$ is not a flat. If the resulting polymatroid $P^{\prime}$ is simple, then take $P^{\mathrm{sim}}:=P^{\prime}$; otherwise, repeat these steps, replacing $(P,\mathbf{n})$ with $(P^{\prime},\mathbf{n}^{\prime})$. This process terminates because the cage shrinks at each step. If the cage is $\mathbf{0}$, then $P$ is the polymatroid of rank 0 on zero elements, which is simple; otherwise, the process stopped before the cage was equal to $\mathbf{0}$, meaning that a simple polymatroid was produced.
Lemmas 3.15 and 3.16 show that $\mathcal{L}_{P^{\mathrm{sim}}}\cong\mathcal{L}_{P}$, and 3.14 implies $P^{\mathrm{sim}}$ is the unique simple polymatroid with this property. This completes the proof of 1.1. ∎
@@
If $P$ is loopless and realized by a subspace arrangement $V_{1},\ldots,V_{N}\subset V$, then $P$ fails to be simple if and only if $V_{i}\subset V_{j}\subsetneq V$ for some $i\neq j$. In this case, a step of the “simplification process” used to construct $P^{\mathrm{sim}}$ in the proof of 1.1 corresponds to replacing $V_{i}$ with $V_{i}+\ell$, where $\ell\subset V$ is a generic line, then deleting loops. Alternatively, one may view this as replacing $V_{i}$ with its immediate successor $V_{i,\operatorname{codim}V_{i}-1}$ in a general flag containing $V_{i}$ (see 3.7 and 3.18).
@@
[…]
@@
We axiomatically characterize posets of combinatorial flats. In light of 1.1, this provides a cryptomorphic definition of simple polymatroids, generalizing the equivalence of geometric lattices and simple matroids.
@@
where $\operatorname{rk}_{\mathcal{L}}(e)$ is the length of any saturated chain in $[\hat{0},e]$. The geometric part of $e$, denoted $e^{\mathrm{geo}}$, is the join of all elements of $[\hat{0},e]$ that are maximal irreducibles of $\mathcal{L}$.
@@
If $P$ is a polymatroid (which we may assume simple by 1.1ii), then the prescribed properties hold for $\mathcal{L}_{P}$ by 1.1i, 3.13, and 3.2.
@@
Since the irreducibles of $\mathcal{L}$ are downward closed, for each $i\in E$, $[\hat{0},i]=\{\hat{0}<x_{i,1}<\cdots<x_{i,n_{i}}=i\}$. By construction of $P$, the following are equivalent:
@@
Plainly, the compositions $\varphi\circ\varphi^{\prime}$ and $\varphi^{\prime}\circ\varphi$ are increasing maps (i.e. $(\varphi\circ\varphi^{\prime})(x)\geq x$ and $(\varphi^{\prime}\circ\varphi)(\mathbf{s})\geq\mathbf{s}$). To finish, we show that both $\varphi$ and $\varphi^{\prime}$ decrease rank. Let $\mathbf{s}$ be a combinatorial flat and let $F=\{i:s_{i}=n_{i}\}$, so that $\mathbf{s}^{\mathrm{geo}}=\sum_{i\in F}n_{i}\mathbf{e}_{i}$. Using submodularity of $\operatorname{rk}_{\mathcal{L}}$, our discussion of flats of $P$, and 3.2, we obtain
@@ 4. Schubert varieties of subspace arrangements
We introduce and study the Schubert variety $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ associated to a linear subspace $V\subset\mathbb{K}^{\mathbf{n}}$, with $\mathbb{K}$ an algebraically closed field of characteristic 0. Our construction is founded on work of Hassett and Tschinkel [HT99], which we review in Sections 4.1 and 4.2. We construct $Y_{V}$ in Section 4.4, then in Section 4.5 compute the support of its class in the Chow ring of $\mathbb{P}^{\mathbf{n}}$. In Section 4.6, we discuss “polymatroid generality”, a condition that guarantees $Y_{V}$ consists of finitely many $V$-orbits. We assume this condition in Section 4.7, where we use the results of Section 4.5 to identify strata of $Y_{V}$, ending with a proof of 1.3. Finally, in Section 4.8, we give a presentation for the singular cohomology ring of $Y_{V}$ when $\mathbb{K}=\mathbb{C}$ and $V$ is polymatroid general.
@@
in which the image of $\lambda_{n}$ normalizes the image of $\rho_{n}$. The orbits of $\mathbb{G}_{a}^{n}\rtimes\mathbb{G}_{m}$ on $\mathbb{P}^{n}$ are the same as those of $\mathbb{G}_{a}^{n}$, but each $\mathbb{G}_{a}^{n}$-orbit $O_{i}$ contains a unique $\mathbb{G}_{m}$-fixed point. In coordinates, the fixed point is $[\mathbf{e}_{n+1-i}]$. Consequently, we may canonically identify $O_{i}$ with $\mathbb{G}_{a}^{n}/\operatorname{St}_{i}$ via the map $\mathbf{a}\mapsto\rho_{n}(\mathbf{a})\cdot[\mathbf{e}_{n+1-i}]$.
@@
Throughout the remainder of Section 4, fix notation:
@@
Phrases like “combinatorial flat of $V$” should be taken to mean “combinatorial flat of $(P,\mathbf{n})$.” In Sections 4.7 and 4.8, we will assume that $V$ is “polymatroid general”, in the sense defined in Section 4.6.
4.4.
We construct Schubert varieties of subspace arrangements, bootstrapping from the Hassett-Tschinkel action of Section 4.1. Identifying $\mathbb{G}_{a}^{n_{i}}$ and $\mathbb{K}^{n_{i}}$, the Hassett-Tschinkel maps $\rho_{n_{i}}:\mathbb{K}^{n_{i}}\to\operatorname{Aut}(\mathbb{P}^{n_{i}})$ assemble to a map $\rho:\mathbb{K}^{\mathbf{n}}\to\operatorname{Aut}(\mathbb{P}^{\mathbf{n}})$, allowing $\mathbb{K}^{\mathbf{n}}$ to act on $\mathbb{P}^{\mathbf{n}}$. For each multiset $\mathbf{s}\leq\mathbf{n}$, there is an $|\mathbf{s}|$-dimensional orbit $O_{\mathbf{s}}:=O_{s_{1}}\times\cdots\times O_{s_{N}}\subset\mathbb{P}^{\mathbf{n}}$, whose closure is a product of projectivized coordinate subspaces $\mathbb{P}^{\mathbf{s}}:=\mathbb{P}^{s_{1}}\times\cdots\times\mathbb{P}^{s_{N}}\subset\mathbb{P}^{\mathbf{n}}$. All points in $O_{\mathbf{s}}$ have the same stabilizer, denoted by $\operatorname{St}_{\mathbf{s}}$. The explicit description in Section 4.1 shows that $\operatorname{St}_{\mathbf{s}}$ is the coordinate subspace of $\mathbb{K}^{\mathbf{n}}$ cut out by
@@
By Section 4.2, $\rho$ extends to an injection
@@
We compute the support of the Chow class of $Y_{V}$. Given a nonsingular variety $X$, let $A_{d}(X)$ be the Chow group of $X$ spanned by $d$-dimensional cycles. If $Y$ and $Y^{\prime}$ are subvarieties of $X$, then their intersection product is written $[Y]\cdot[Y^{\prime}]$.
@@
$[Y_{V}]=\sum_{\mathbf{b}\text{ basis of $P$}}c_{\mathbf{b}}\,[\mathbb{P}^{\mathbf{b}}].$
@@
Write $[Y_{V}]=\sum_{|\mathbf{t}|=d}c_{\mathbf{t}}[\mathbb{P}^{\mathbf{t}}]$. Let $\mathbf{s}$ be a multiset with $|\mathbf{s}|=d$ that is not a basis of $P$. We will show that $[\mathbb{P}^{\mathbf{s}}]$ has coefficient zero in $[Y_{V}]$.
Let $T=\{i:\overline{s}_{i}=n_{i}\}$. This set is nonempty: if we suppose falsely that $T=\emptyset$, then by 3.2, $\mathbf{s}=\overline{\mathbf{s}}$. This means that $\mathbf{s}$ is a combinatorial flat satisfying $\mathbf{s}^{\mathrm{geo}}=\mathbf{0}$, so by 3.2 once more,
@@
The class $[\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot[Y_{V}]$ is represented by a subvariety $Y^{\prime}$ of $Y_{V}$, obtained by intersecting $Y_{V}$ with appropriately general hyperplanes. To show that $[\mathbb{P}^{\pi_{T}(\mathbf{n}-\mathbf{s}^{\prime})}]\cdot\pi_{T*}([\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot[Y_{V}])=[\mathbb{P}^{\pi_{T}(\mathbf{n}-\mathbf{s}^{\prime})}]\cdot\pi_{T*}([Y^{\prime}])$ is zero, it suffices to show that $\dim\pi_{T}(Y^{\prime})<|\pi_{T}(\mathbf{s}^{\prime})|$. In fact, this is the case: by 3.2,
@@
$[\overline{V}]=\sum_{\text{$\mathbf{b}$ basis of $P$}}[\mathbb{P}^{\mathbf{b}}].$
Therefore if we can find a flat family $X_{t}$ of subvarieties of $\mathbb{P}^{\mathbf{n}}$ such that $X_{1}$ is isomorphic to $Y_{V}$ and $X_{0}\supset\overline{V}$, then $c_{\mathbf{b}}\geq 1$ for all bases $\mathbf{b}$ of $P$. Recall from Section 4.4 that $Y_{V}$ is the closure of $\iota(V)$, where $\iota:\mathbb{K}^{\mathbf{n}}\to O_{\mathbf{n}}$. The idea is that $I(V)$ is the tangent space to $\iota(V)$ at the origin of $O_{\mathbf{n}}$, so we can construct a deformation of $Y_{V}$ by taking the closure of the deformation of $\iota(V)$ to its tangent space.
@@
From Section 4.1, we see that
@@
Another proof of 4.6 goes as follows: a projection of $O_{\mathbf{n}}\cap Y_{V}$ onto a given $d$-element subset of the coordinates $\{b_{ij}/b_{i0}\}_{i,j}$ is dominant if and only if its induced map on tangent spaces is surjective at a general point of $O_{\mathbf{n}}\cap Y_{V}$. Consequently, to prove that $c_{\mathbf{b}}>0$, it suffices to produce a point $p\in O_{\mathbf{n}}\cap Y_{V}$ and a $d$-element set of coordinates representing $\mathbf{b}$ such that the derivative of the corresponding projection is surjective at $p$.
@@
If the matroid of $V$ in $(\mathbb{K}^{1})^{|\mathbf{n}|}\cong\mathbb{K}^{\mathbf{n}}$ is $\widetilde{P}$, then $V$ is polymatroid general, so any realizable polymatroid has a p.g. realization by 2.10. However, not all p.g. realizations of $P$ realize $\widetilde{P}$.

### v3

@@ 3. Combinatorial flats
By the matroid case and ˜2.15, a set $S$ is a flat of $\widetilde{T_{F}P}$ if and only if $S$ is a flat of $\widetilde{P}$ that either contains $\cup_{i\in F}E_{i}$ or satisfies $\overline{S\cup i}\not\supset\cup_{i\in F}E_{i}$ for all $i\in\widetilde{E}\setminus S$. In the former case, $\operatorname{rk}_{\widetilde{T_{F}P}}(S)=\operatorname{rk}_{\widetilde{P}}(S)-1$; in the latter $\operatorname{rk}_{\widetilde{T_{F}P}}(S)=\operatorname{rk}_{\widetilde{P}}(S)$. Taking $\mathbf{s}$ to be the multiset represented by $S$ completes the proof. ∎
In the remainder of Section˜3, we work up to proving ˜1.1, then provide axioms for combinatorial flats. These results are not required to understand Schubert varieties of subspace arrangements; readers interested solely in geometry may proceed to Section˜4 forthwith.
$3.$
@@
Semimodularity is ˜3.11.
[…]
@@
[…]
@@
[…]
@@
Generalizations of Lemmas 3.11 and 3.12 for matricubes appear in [AG24, Theorem 3.4] (see ˜3.8). With Lemmas 3.11 and 3.12 in hand, we have established ˜1.1 parts (i), (iii), and (iv). We now devote our attention to ˜1.1(ii), which states that $P$ is determined by $\mathcal{L}_{P}$ up to simplification.
@@
Since $P$ is simple, $E_{i}$ is a flat of $\widetilde{P}$ for all $i\in E$. The restriction of $\widetilde{P}$ to $E_{i}$ is a boolean matroid of rank $n_{i}=\operatorname{rk}_{P}(i)$, so all proposed join-irreducibles are combinatorial flats of $(P,\mathbf{n})$. Each $s_{i}\mathbf{e}_{i}$ is join-irreducible because the interval $\mathbf{0},s_{i}\mathbf{e}_{i}]\subset\mathcal{L}_{P,\mathbf{n}}$ is a chain of length $s_{i}$. No other combinatorial flats are join-irreducible because $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$ can be written $\mathbf{s}=s_{1}\mathbf{e}_{1}\vee\cdots\vee s_{N}\mathbf{e}_{N}$. ∎
@@
By ˜3.13, each $i\in E$ corresponds to a maximal join-irreducible $x_{i}$ of $\mathcal{L}_{P,\mathbf{n}}$. The rank function of $P$ is recovered by $\operatorname{rk}_{P}(S)=\operatorname{rk}_{P,\mathbf{n}}(\vee_{i\in S}x_{i})$. ∎
@@
We show that reduction and loops do not change the lattice of combinatorial flats. From these facts, we derive ˜1.1.
@@
For surjectivity, suppose that $\mathbf{s}\in\mathcal{L}_{R_{i}(P,\mathbf{n})}$ is represented by a flat $S$ of $\widetilde{R_{i}(P,\mathbf{n})}$. We will show there is $\mathbf{s}^{\prime}\in\mathcal{L}_{P,\mathbf{n}}$ with $\psi(\mathbf{s}^{\prime})=\mathbf{s}$. By Lemmas 2.17 and 2.12, there is $e\in E_{i}\setminus S$ such that $S\cup e$ or $S$ is a flat of $\widetilde{P}$. If $S$ is a flat of $\widetilde{P}$, then $s_{i}<n_{i}$, so we may take $\mathbf{s}^{\prime}=\mathbf{s}$. Otherwise, suppose $S\cup e$ is a flat of $\widetilde{P}$. If $s_{i}+1<n_{i}$, then by ˜3.2, $\mathbf{s}$ is a combinatorial flat of $\widetilde{P}$, so we may again take $\mathbf{s}^{\prime}=\mathbf{s}$. Otherwise, if $S\cup e$ is a flat of $\widetilde{P}$ and $s_{i}+1=n_{i}$, then take $\mathbf{s}^{\prime}=\mathbf{s}+\mathbf{e}_{i}$. This completes the proof of surjectivity.
@@
Proof of ˜1.1.
Define $\mathcal{L}_{P}:=\mathcal{L}_{P,\mathbf{n}}$, with $\mathbf{n}$ any cage for $P$. This definition does not depend on the choice of cage because we may assume that $P$ is loopless by ˜3.16, then repeatedly apply ˜3.15 to produce an isomorphism $\mathcal{L}_{P,\mathbf{n}}\cong\mathcal{L}_{P,(\operatorname{rk}_{P}(1),\ldots,\operatorname{rk}_{P}(N))}$. We will henceforth assume that $\mathbf{n}=(\operatorname{rk}_{P}(1),\ldots,\operatorname{rk}_{P}(N))$.
Statements ˜i and ˜iv are the content of ˜3.12; we now move to proving ˜ii. To construct $P^{\mathrm{sim}}$: let $(P^{\prime},\mathbf{n}^{\prime})=R_{i}(P\setminus L,\pi_{E\setminus L}(\mathbf{n}))$, where $L$ is the set of loops of $P$ and $i\in E\setminus L$ is any element such that $\{i\}$ is not a flat. If the resulting polymatroid $P^{\prime}$ is simple, then take $P^{\mathrm{sim}}:=P^{\prime}$; otherwise, repeat these steps, replacing $(P,\mathbf{n})$ with $(P^{\prime},\mathbf{n}^{\prime})$. This process terminates because the cage shrinks at each step. If the cage is $\mathbf{0}$, then $P$ is the polymatroid of rank 0 on zero elements, which is simple; otherwise, the process stopped before the cage was equal to $\mathbf{0}$, meaning that a simple polymatroid was produced.
Lemmas 3.15 and 3.16 show that $\mathcal{L}_{P^{\mathrm{sim}}}\cong\mathcal{L}_{P}$, and ˜3.14 implies $P^{\mathrm{sim}}$ is the unique simple polymatroid with this property. This completes the proof of ˜1.1. ∎
@@
If $P$ is loopless and realized by a subspace arrangement $V_{1},\ldots,V_{N}\subset V$, then $P$ fails to be simple if and only if $V_{i}\subset V_{j}\subsetneq V$ for some $i\neq j$. In this case, a step of the “simplification process” used to construct $P^{\mathrm{sim}}$ in the proof of ˜1.1 corresponds to replacing $V_{i}$ with $V_{i}+\ell$, where $\ell\subset V$ is a generic line, then deleting loops. Alternatively, one may view this as replacing $V_{i}$ with its immediate successor $V_{i,\operatorname{codim}V_{i}-1}$ in a general flag containing $V_{i}$ (see ˜3.7 and ˜3.18).
@@
$\mathcal{L}_{P,\mathbf{n}}=\{\mathbf{0},\mathbf{e_{1}},\mathbf{e_{2}},2\mathbf{e_{1}}+2\mathbf{e_{2}}\}$$V_{1}=V_{2}$$V_{1,1}$$V_{2,1}$
$\mathcal{L}_{R_{1}(P,\mathbf{n})}=\{\mathbf{0},\mathbf{e_{1}},\mathbf{e_{2}},\mathbf{e_{1}}+2\mathbf{e_{2}}\}$$V_{2}$$V_{1}$$V_{2,1}$
$\mathcal{L}_{R_{2}(R_{1}(P,\mathbf{n}))}=\{\mathbf{0},\mathbf{e_{1}},\mathbf{e_{2}},\mathbf{e_{1}}+\mathbf{e_{2}}\}$$V_{1}$$V_{2}$
@@
We axiomatically characterize posets of combinatorial flats. In light of ˜1.1, this provides a cryptomorphic definition of simple polymatroids, generalizing the equivalence of geometric lattices and simple matroids.
@@
where $\operatorname{rk}_{\mathcal{L}}(e)$ is the length of any saturated chain in $\hat{0},e]$. The geometric part of $e$, denoted $e^{\mathrm{geo}}$, is the join of all elements of $\hat{0},e]$ that are maximal irreducibles of $\mathcal{L}$.
@@
If $P$ is a polymatroid (which we may assume simple by ˜1.1˜ii), then the prescribed properties hold for $\mathcal{L}_{P}$ by ˜1.1˜i, ˜3.13, and ˜3.2.
@@
Since the irreducibles of $\mathcal{L}$ are downward closed, for each $i\in E$, $\hat{0},i]=\{\hat{0}<x_{i,1}<\cdots<x_{i,n_{i}}=i\}$. By construction of $P$, the following are equivalent:
@@
Plainly, the compositions $\varphi\circ\varphi^{\prime}$ and $\varphi^{\prime}\circ\varphi$ are increasing maps (i.e. $(\varphi\circ\varphi^{\prime})(x)\geq x$ and $(\varphi^{\prime}\circ\varphi)(\mathbf{s})\geq\mathbf{s}$). To finish, we show that both $\varphi$ and $\varphi^{\prime}$ decrease rank. Let $\mathbf{s}$ be a combinatorial flat and let $F=\{i:s_{i}=n_{i}\}$, so that $\mathbf{s}^{\mathrm{geo}}=\sum_{i\in F}n_{i}\mathbf{e}_{i}$. Using submodularity of $\operatorname{rk}_{\mathcal{L}}$, our discussion of flats of $P$, and ˜3.2, we obtain
@@ 4. Schubert varieties of subspace arrangements
We introduce and study the Schubert variety $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ associated to a linear subspace $V\subset\mathbb{K}^{\mathbf{n}}$, with $\mathbb{K}$ an algebraically closed field of characteristic 0. Our construction is founded on work of Hassett and Tschinkel [HT99], which we review in Sections 4.1 and 4.2. We construct $Y_{V}$ in Section˜4.4, then in Section˜4.5 compute the support of its class in the Chow ring of $\mathbb{P}^{\mathbf{n}}$. In Section˜4.6, we discuss “polymatroid generality”, a condition that guarantees $Y_{V}$ consists of finitely many $V$-orbits. We assume this condition in Section˜4.7, where we use the results of Section˜4.5 to identify strata of $Y_{V}$, ending with a proof of ˜1.3. Finally, in Section˜4.8, we give a presentation for the singular cohomology ring of $Y_{V}$ when $\mathbb{K}=\mathbb{C}$ and $V$ is polymatroid general.
@@
in which the image of $\lambda_{n}$ normalizes the image of $\rho_{n}$. The orbits of $\mathbb{G}_{a}^{n}\rtimes\mathbb{G}_{m}$ on $\mathbb{P}^{n}$ are the same as those of $\mathbb{G}_{a}^{n}$, but each $\mathbb{G}_{a}^{n}$-orbit $O_{i}$ contains a unique $\mathbb{G}_{m}$-fixed point. In coordinates, the fixed point is $\mathbf{e}_{n+1-i}]$. Consequently, we may canonically identify $O_{i}$ with $\mathbb{G}_{a}^{n}/\operatorname{St}_{i}$ via the map $\mathbf{a}\mapsto\rho_{n}(\mathbf{a})\cdot[\mathbf{e}_{n+1-i}]$.
@@
Throughout the remainder of Section˜4, fix notation:
@@
Phrases like “combinatorial flat of VV” should be taken to mean “combinatorial flat of $(P,\mathbf{n})$.” In Sections 4.7 and 4.8, we will assume that $V$ is “polymatroid general”, in the sense defined in Section˜4.6.
$4.$
We construct Schubert varieties of subspace arrangements, bootstrapping from the Hassett-Tschinkel action of Section˜4.1. Identifying $\mathbb{G}_{a}^{n_{i}}$ and $\mathbb{K}^{n_{i}}$, the Hassett-Tschinkel maps $\rho_{n_{i}}:\mathbb{K}^{n_{i}}\to\operatorname{Aut}(\mathbb{P}^{n_{i}})$ assemble to a map $\rho:\mathbb{K}^{\mathbf{n}}\to\operatorname{Aut}(\mathbb{P}^{\mathbf{n}})$, allowing $\mathbb{K}^{\mathbf{n}}$ to act on $\mathbb{P}^{\mathbf{n}}$. For each multiset $\mathbf{s}\leq\mathbf{n}$, there is an $|\mathbf{s}|$-dimensional orbit $O_{\mathbf{s}}:=O_{s_{1}}\times\cdots\times O_{s_{N}}\subset\mathbb{P}^{\mathbf{n}}$, whose closure is a product of projectivized coordinate subspaces $\mathbb{P}^{\mathbf{s}}:=\mathbb{P}^{s_{1}}\times\cdots\times\mathbb{P}^{s_{N}}\subset\mathbb{P}^{\mathbf{n}}$. All points in $O_{\mathbf{s}}$ have the same stabilizer, denoted by $\operatorname{St}_{\mathbf{s}}$. The explicit description in Section˜4.1 shows that $\operatorname{St}_{\mathbf{s}}$ is the coordinate subspace of $\mathbb{K}^{\mathbf{n}}$ cut out by
@@
By Section˜4.2, $\rho$ extends to an injection
@@
We compute the support of the Chow class of $Y_{V}$. Given a nonsingular variety $X$, let $A_{d}(X)$ be the Chow group of $X$ spanned by $d$-dimensional cycles. If $Y$ and $Y^{\prime}$ are subvarieties of $X$, then their intersection product is written $Y]\cdot[Y^{\prime}]$.
@@
$Y_{V}]=\sum_{\mathbf{b}\text{ basis of $P$}}c_{\mathbf{b}}\,[\mathbb{P}^{\mathbf{b}}].$
@@
Write $Y_{V}]=\sum_{|\mathbf{t}|=d}c_{\mathbf{t}}[\mathbb{P}^{\mathbf{t}}]$. Let $\mathbf{s}$ be a multiset with $|\mathbf{s}|=d$ that is not a basis of $P$. We will show that $\mathbb{P}^{\mathbf{s}}]$ has coefficient zero in $Y_{V}]$.
Let $T=\{i:\overline{s}_{i}=n_{i}\}$. This set is nonempty: if we suppose falsely that $T=\emptyset$, then by ˜3.2, $\mathbf{s}=\overline{\mathbf{s}}$. This means that $\mathbf{s}$ is a combinatorial flat satisfying $\mathbf{s}^{\mathrm{geo}}=\mathbf{0}$, so by ˜3.2 once more,
@@
The class $\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot[Y_{V}]$ is represented by a subvariety $Y^{\prime}$ of $Y_{V}$, obtained by intersecting $Y_{V}$ with appropriately general hyperplanes. To show that $\mathbb{P}^{\pi_{T}(\mathbf{n}-\mathbf{s}^{\prime})}]\cdot\pi_{T*}([\mathbb{P}^{\mathbf{n}-\mathbf{s}+\mathbf{s}^{\prime}}]\cdot[Y_{V}])=[\mathbb{P}^{\pi_{T}(\mathbf{n}-\mathbf{s}^{\prime})}]\cdot\pi_{T*}([Y^{\prime}])$ is zero, it suffices to show that $\dim\pi_{T}(Y^{\prime})<|\pi_{T}(\mathbf{s}^{\prime})|$. In fact, this is the case: by ˜3.2,
@@
[…]
Therefore if we can find a flat family $X_{t}$ of subvarieties of $\mathbb{P}^{\mathbf{n}}$ such that $X_{1}$ is isomorphic to $Y_{V}$ and $X_{0}\supset\overline{V}$, then $c_{\mathbf{b}}\geq 1$ for all bases $\mathbf{b}$ of $P$. Recall from Section˜4.4 that $Y_{V}$ is the closure of $\iota(V)$, where $\iota:\mathbb{K}^{\mathbf{n}}\to O_{\mathbf{n}}$. The idea is that $I(V)$ is the tangent space to $\iota(V)$ at the origin of $O_{\mathbf{n}}$, so we can construct a deformation of $Y_{V}$ by taking the closure of the deformation of $\iota(V)$ to its tangent space.
@@
From Section˜4.1, we see that
@@
Another proof of ˜4.6 goes as follows: a projection of $O_{\mathbf{n}}\cap Y_{V}$ onto a given $d$-element subset of the coordinates $\{b_{ij}/b_{i0}\}_{i,j}$ is dominant if and only if its induced map on tangent spaces is surjective at a general point of $O_{\mathbf{n}}\cap Y_{V}$. Consequently, to prove that $c_{\mathbf{b}}>0$, it suffices to produce a point $p\in O_{\mathbf{n}}\cap Y_{V}$ and a $d$-element set of coordinates representing $\mathbf{b}$ such that the derivative of the corresponding projection is surjective at $p$.
@@
If the matroid of $V$ in $(\mathbb{K}^{1})^{|\mathbf{n}|}\cong\mathbb{K}^{\mathbf{n}}$ is $\widetilde{P}$, then $V$ is polymatroid general, so any realizable polymatroid has a p.g. realization by ˜2.10. However, not all p.g. realizations of $P$ realize $\widetilde{P}$.